Ethanol, C2H5OH, is also highly flammable and combusts according to the reaction:


C2H5OH (l) + 3 O2 (g) → 3 H2O (g) + 2 CO2 (g) + 1270 kJ


a.)If 67.9 grams of ethanol combusts inside a sealed container with 102 grams of O2, how much heat will be released by the chemical reaction?


b.)Assume all of the heat released in (a) is absorbed by the 840 grams of air surrounding the reaction vessel, originally at 16 Celsius. Assuming the air absorbed all the heat, determine the final temperature of the air. The specific heat of air is 1.005 J/gC.


1
Expert's answer
2021-03-10T06:54:08-0500

moles of ethanol = 67.9/46 = 1.47

moles of oxygen = 102/32 = 3.18

3 moles of oxygen = 1270 KJ

3.18 moles of oxygen = 3.18 ×(1270/3) = 1349.3 KJ


Q = mc "\Delta T"

1349300 J = 840 × 1.005 × "\Delta T"

"\Delta T" = 1678.23

"\Delta T" = T2 - T1

1678.22 = T2 --16

T2 = 1694.23⁰C